OpenStudy (anonymous):

At equilibrium, what is the hydrogen-ion concentration if the acid dissociation constant is 0.000 001 and the acid concentration is 0.01M?

OpenStudy (aaronq):

HA <-> H+ + A- Ka=[H+][A-]/[HA] [H+]=[A-]=x Ka=x^2/[HA]

OpenStudy (anonymous):

what does H stand for?

OpenStudy (anonymous):

or is it the answer?

OpenStudy (aaronq):

[H+] = would be be hydrogen cation concentration

OpenStudy (anonymous):

oh A stands for Acid?

OpenStudy (aaronq):

which is what you are looking for, just plug in your values into that last equation and solve for x

OpenStudy (anonymous):

can you help me with that?

OpenStudy (aaronq):

HA is any acid, A- is the conjugate base

OpenStudy (aaronq):

Ka=x^2/[HA] Ka=0.000 001 [HA]= 0.01M

OpenStudy (anonymous):

ok itd be like 0.000 001= x^2/ 0.1?

OpenStudy (anonymous):

@Festinger

OpenStudy (festinger):

yes.

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

do I move 0.1 to 0.000 001 and add them?

OpenStudy (festinger):

\[ 0.000 001= \frac{x^{2}}{0.1}\] \[ 0.000 001*0.1 = x^{2}\]

OpenStudy (anonymous):

ok itd be 0.00000001^2?

OpenStudy (anonymous):

@Festinger

OpenStudy (festinger):

You need to brush up on your algebra. \[0.000 001*0.1 = x^{2}\]\[0.000 000 1 = x^{2}\]\[x=\sqrt{0. 000 000 1}\] pH is defined as -log[H]. Thus: \[-log[H]=-log x = 3.5\]